Google Sheets can be used to find the critical value for an F distribution for a given significance level. For example, to find the critical value (the value above which you would reject the null hypothesis) for α=0.05α=0.05 with dfbetween=3dfbetween=3 and dfwithin=24dfwithin=24, enter

=FINV(0.05,3,24)

Enter this into Google Sheets to confirm you obtain the value 3.009.



You conduct a one-factor ANOVA with 6 groups and 5 subjects in each group (a balanced design). Use Google Sheets to find the critical values for α=0.1α=0.1 and α=0.05α=0.05 (report accurate to 3 decimal places).
F0.1=F0.1=
F0.05=F0.05=
Answer #1

here dfbetween= number of groups k-1= 6-1 =5

dfwithin= nk-n=6*5-6=24

therefore

critical values for α=0.1 =finv(0.1,5,24)=2.103

critical values for α=0.05 =finv(0.05,5,24)=2.621